OpenStudy (anonymous):

What is the value of Kw at 25°C?

thomaster (thomaster):

\(\Large K_w=\dfrac{[H^+][OH^-]}{[H_2O]}\) Only 1 molecule in 10 million will be dissociated. so H+ and OH- are both \(\large10^{-7}\) \(\Large K_w=\dfrac{[10^{-7}][10^{-7}]}{[1]}=1.0*10^{-14}\)

OpenStudy (anonymous):

Thanks!

thomaster (thomaster):

No problem :)